# Matching Hadronization and Perturbative Evolution: The Cluster Model in Light of Infrared Shower Cutoff Dependence

### v1

@@ 7.2 Tuning Quality and Cutoff Dependence of the Tuning Parameters
At this point an interesting aspect to be discussed is to which extent the reference data used for the tuning analyses, which are generated by the MC for the cutoff value $Q_{0,\mathrm{ref}}=1.25$ GeV, represent a good proxy for the real LEP data. Furthermore, it should be also clarified whether the experimental data indeed prefers shower cut values in the range where perturbation theory is still valid, i.e. in the range between $1$ and $2$ GeV that we consider in our discussion. To address these questions we carry out two additional tuning analyses where instead of the MC generated reference data the LEP data is used that also enters the regular Herwig release tunes. In one tuning analysis $Q_{0}$ is treated as a freely floating tuning parameter and in the other the shower cutoff is fixed to $Q_{0,\mathrm{ref}}=1.25$ GeV. The outcome is shown as the colored triangles for the cubic as well as for the quartic interpolations. The minimal GOF function values for these tuning fits are of course not zero since the real data always differs from simulations. For the cubic interpolation both tunes yields practically identical values, with minimal GOF values compatible with the smaller dynamical model GOF values for $Q_{0}>1.4$ GeV. For the quartic interpolation the minimal GOF values for the tune with a freely floating shower cutoff yields $Q_{0}$ values close to unity, and the minimal GOF values are in the range from $4.2$ to $4.6$ for the default and from $3.57$ to $3.64$ for the dynamical model. The results confirm that the MC reference data of our shower cutoff dependent tuning analysis is sufficiently close to the experimental LEP data and, more importantly, that the interval of $1$ to $2$ GeV, where the shower cut $Q_{0}$ can be considered as an infrared factorization scale, is well compatible with the experimental data.
@@
Let is now have a look at the $Q_{0}$-dependence of the six hadronization model parameters which are treated as floating parameters in our tuning analyses in each of the hadronization models. As we have already mentioned in Sec. 6.2, for the dynamical model a proper matching to the parton shower (which allows for the shower cutoff $Q_{0}$ to be interpreted as an infrared factorization scale) implies a linear $Q_{0}$-dependence for the “hard” starting scales $\tilde{Q}_{g}$ and $\tilde{Q}_{q}$ of the non-perturbative branching processes and an insensitivity to $Q_{0}$ for the other hadronization parameters. On the other hand, for the default hadronization model (where the shower cutoff is merely another hadronization parameter) similar implications can in principle not necessarily be expected. In Figs. 8 and 9 we show the $Q_{0}$-dependence of the six tuning parameters which were treated as floating parameters for the default and the dynamic models, respectively. We show the results based on the cubic (orange) and the quartic (blue) interpolations. The dots show the central values ${p_{i,\mathrm{cent}}(Q_{0})}$ and the vertical lines stand for the uncertainties which are obtained from the quadratic sum of the statistical und interpolation uncertainty $\Delta p_{i}=\sqrt{(\Delta p_{i,\mathrm{stat}}(Q_{0}))^{2}+(\Delta p_{i,\mathrm{inter}}(Q_{0}))^{2}}$, as described in Sec. 6.2. Overall we see that that tuning results based on the cubic and quartic interpolations are nicely compatible, even though the error bars do not always overlap.
@@
In Fig. 9 we see that, indeed, the parameters $\tilde{Q}_{g}$ and $\tilde{Q}_{q}$ exhibit a $Q_{0}$-dependence that is quite close to linear. At the same time, in comparison the other four hadronization parameters (${\mathtt{Cl}_{\mathtt{max}}}$, ${\mathtt{Cl}_{\mathtt{pow}}}$, $\mathtt{PwtSquark}$ and $\mathtt{PwtDIquark}$) are, indeed, rather $Q_{0}$-insensitive. The behavior is again not perfect, with ${\mathtt{Cl}_{\mathtt{max}}}$ and ${\mathtt{Cl}_{\mathtt{pow}}}$ exhibiting rather large uncertainties, but it can nevertheless be observed rather clearly. In contrast, we can see a somewhat larger $Q_{0}$-dependence of the same four parameters for the default model in Fig. 8. This is visible most prominently for $\mathtt{PwtDIquark}$. Since these parameters directly affect the Baryon production and thus the charge particle multiplicities generated in the simulation, we can conclude that in the default model the shower cutoff value $Q_{0}$ affects these multiplicities so that the tuned value of $\mathtt{PwtDIquark}$ needs to compensate in a significant way. In the novel dynamical hadronization model this rather unnatural feature is not visible. Overall, the results shown in Figs. 8 and 9 again confirm that the novel dynamical hadronization model achieves a better separation of the low-scale hadronization processes from the parton level description provided by the parton shower. This separation is essentially the MC simulation analogue of factorization in analytic QCD studies and an important prerequisite for the parton shower cutoff to be interpreted as a factorization scale.
@@ 7.3 Rescaled Migration Matrix Function
Before moving on, let us comment on a general feature of the migration functions. It concerns that $S^{\rm MC}(\ell,\{\hat{k},Q,Q_{0}\})$ is finite for $\ell$ values down to $-1$ GeV. We can see this feature clearly in the right panels of Figs. 10 and Fig. 11 for partonic values $\hat{k}>1$ GeV. Even though the major portion of $S^{\rm MC}(\ell,\{\hat{k},Q,Q_{0}\})$ is located at positiv $\ell$, which means that the bulk of hadronization corrections shift the thrust distribution towards larger thrust values, this negative tail shows that hadronization can sometimes also decrease the thrust value. Since this feature arises from the tuning to actual LEP data, it means that at least for the Herwig simulation implementation this particular feature of the hadronization corrections is demanded by data. For $\hat{k}\geq 1.5$ GeV this behaviour arises consistently and becomes independent of the value of $\hat{k}$ also for other $Q$ and $Q_{0}$. This feature is per se not at all problematic and also consistent with a shape function within QCD factorization. However, for partonic $\hat{k}$ values below $1.5$ GeV the migration function is not capable to build up such a negative tail, simply because the physical thrust values are restricted to be positive. This entails that the first $\ell$-moment of the rescaled migration function $S^{\rm MC}(\ell,\{\hat{k},Q,Q_{0}\})$ (see Eq. (13)) for $\hat{k}$ values below $1.5$ GeV is always larger than for values larger than $2$ GeV, where the first moment stabilizes. This feature is absent in the shape function $S_{\rm had}(\ell,Q_{0})$, which is strictly $\hat{k}$-independent. We come back to this feature in our discussion below.
@@ 7.4 Shower Cutoff and Energy Dependence of the Migration Matrix Function
where $\Delta_{\rm soft}(Q_{0},Q_{0,\mathrm{ref}})$ is determined from the $R$-evolution equation in Eq. (10).
@@
It is instructive to discuss this failure of the default model from a the perspective of overall size of the hadronization corrections. As we can see from Fig. 14, for $Q_{0,\mathrm{ref}}=1.25$ we have $\Omega_{1}^{\rm MC}$ in the range between $1.2$ and $1.5$ GeV for partonic momenta $\hat{k}>2$ GeV. This should be compared to the overall variation of $\Omega_{1}^{\rm MC}$ in the range between $Q_{0}=1.0$ GeV and $Q_{0}=1.75$ GeV predicted by $R$-evolution equation (10) obtained from QCD factorization. The latter amounts to around $0.5$ GeV, as we see in the lower panels of Fig. 15. Howerver, because the choice of reference shower cut $Q_{0,\mathrm{ref}}$ is arbitrary, and we do not have any principle means to tell for which value of $Q_{0}$ the default model may have the best agreement with the expectations from QCD factorization (which can only be tested by a $Q_{0}$-dependence consistent with QCD factorization), this implies that the size of the hadronization corrections to the (true) partonic thrust distribution that is provided by the default model is inconsistent with QCD factorization in the range of $40$%. This is substantially worse in comparison to the new default model, where we have consistency at the level of better than $10$%. In the context of the shower cut $Q_{0}$ adopting the role of an IR factorization scale, so that the hadronization effects have a well-defined and controlled scheme dependence so that they can be assigned field theoretic meaning, the default model thus fails quite badly. In this respect the novel dynamical hadronization model performs substantially better.
@@ 8 Conclusions and Outlook
In this article we further promote the idea of the parton shower cutoff $Q_{0}$ for MC simulations being an infrared factorization scale that separates, in a controlled manner and compatible with QCD, perturbative and non-perturbative hadronization effects. In this context, features of the MC’s hadronization model may be given a systematic field theoretic meaning and, at the same time, QCD parameters appearing in the parton shower may be related to their renormalization scheme dependent counter parts appearing in analytic QCD computations. An important prerequisite to achieve this is to have a parton shower algorithms that has at least NLL precision for the observable considered. However, the second essential and equally important prerequisite is to have a hadronization model that can properly match the unavoidable infrared cutoff $Q_{0}$-dependence of the parton level description that emerges from the parton shower. This is a highly nontrivial condition on the hadronization model since it entails that the parton shower cutoff $Q_{0}$ is not treated as a tuned hadronization parameter. Rather, the hadron level MC observable description should be equivalent for different $Q_{0}$ values, at least within some low energy interval, where QCD perturbation theory can still be trusted.
@@
In this work we have investigated these features using the angular ordered parton shower and the cluster hadronization model implemented in the Herwig 7.2 MC event generator focusing primarily on the 2-jettiness distribution in $e^{+}e^{-}$ annihilation for which the angular ordered parton shower is NLL precise. In the earlier work Hoang:2018zrp some of us have analyzed the gluon transverse momentum cutoff $Q_{0}$-dependence that emerges from the angular order parton shower for the 2-jettiness distribution in detail at the subleading ${\cal O}(\alpha_{s})$ level using QCD factorization in the dijet region. It was also shown that the Herwig parton shower implementation satisfies the $Q_{0}$ evolution equation from QCD factorization very well. In this work we have now extended our investigations concerning the parton shower infrared cutoff being a factorization scale from the perspective of the Herwig cluster hadronization model.
@@
The default cluster hadronization model has not been designed in a way so that it systematically matches to the parton shower for a range of $Q_{0}$ values. This is related to a number of features in the cluster formation and fission dynamics that yield good data description through the tuning, but are otherwise ad-hoc and not compatible with the processes that happen in the parton shower evolution close to the cutoff $Q_{0}$. As a result, the hadronization effects to the 2-jettiness distribution provided by the default Herwig cluster model do not satisfy the QCD constraints that emerge from $Q_{0}$ being promoted to a factorization scale in a satisfactory manner. We have demonstrated this in a number of tests based on tuning analyses for different $Q_{0}$ shower cutoff values where $Q_{0}$ is treated as an external scale and not a tuned parameter. These tests involve analyses of the $Q_{0}$-dependence of hadron level simulations for 2-jettiness and also other event-shapes and observables and of the parton-to-hadron migration matrix for 2-jettiness for which QCD factorization provides nontrivial constraints.
@@
To improve the cluster hadronization model we have added a number of features to the cluster formation and cluster fission processes that mimic the gluon emission and splitting dynamics that takes place in the parton shower. These modifications provide a clearer separation of model parameters that are expected to be correlated to $Q_{0}$, governing the ’hard’ aspects of the cluster hadronization dynamics, from those which govern ’soft’ hadron formation aspects that should be rather $Q_{0}$-independent. In our $Q_{0}$-dependent tuning studies we found that this novel dynamical cluster hadronization model performs substantially better concerning the $Q_{0}$-invariance of Herwig’s hadron level predictions as well as concerning the QCD factorization constraints on the 2-jettiness parton-to-hadron migration matrix. We emphasize that our analyses also involved hard scattering energies that are not accounted for in the reference data used for the tuning. This shows that the novel dynamical hadronization model properly scales the consistency to QCD factorization to other hard scattering energies.
@@
Even though the novel dynamical hadronization model we have designed in this article is not perfect, its features provide an important step forward in promoting the hadronization corrections encoded in MC generators to have a well-defined scheme in the QCD context, as also discussed previously in Ref. Platzer:2022jny . This is an essential aspect that should be followed in parallel to the ongoing developments of subleading order precise parton shower algorithms such that the parameters of the parton shower as well as the hadronization model can acquire a systematic QCD field theoretic meaning. Beyond a more precise and consistent description of experimental data, important potential applications of such improvements are the determination of QCD parameters directly from MC studies as well as a systematic quantification of hadronization corrections to analytic QCD calculations from MC simulations. In an upcoming article we will investigate the former application from the perspective of the MC top quark mass parameter.

### v2

@@ 7.2 Tuning Quality and Cutoff Dependence of the Tuning Parameters
At this point an interesting aspect to be discussed is to which extent the reference data used for the tuning analyses, which are generated by the MC for the cutoff value $Q_{0,\mathrm{ref}}=1.25$ GeV, represent a good proxy for the real LEP data. Furthermore, it should be also clarified whether the experimental data indeed prefers shower cut values in the range where perturbation theory is still valid, i.e. in the range between $1$ and $2$ GeV that we consider in our discussion. To address these questions we carry out two additional tuning analyses where instead of the MC generated reference data the LEP data is used that also enters the regular Herwig release tunes. In one tuning analysis $Q_{0}$ is treated as a freely floating tuning parameter and in the other the shower cutoff is fixed to $Q_{0,\mathrm{ref}}=1.25$ GeV. The outcome is shown as the colored triangles for the cubic as well as for the quartic interpolations. The minimal GOF function values for these tuning fits are of course not zero since the real data always differs from simulations. For the cubic interpolation both tunes yield practically identical values, with minimal GOF values compatible with the smaller dynamical model GOF values for $Q_{0}>1.4$ GeV. For the quartic interpolation the minimal GOF values for the tune with a freely floating shower cutoff yields $Q_{0}$ values close to unity, and the minimal GOF values are in the range from $4.2$ to $4.6$ for the default and from $3.57$ to $3.64$ for the dynamical model. The results confirm that the MC reference data of our shower cutoff dependent tuning analysis is sufficiently close to the experimental LEP data and, more importantly, that the interval of $1$ to $2$ GeV, where the shower cut $Q_{0}$ can be considered as an infrared factorization scale, is well compatible with the experimental data.
@@
Let is now have a look at the $Q_{0}$-dependence of the six hadronization model parameters which are treated as floating parameters in our tuning analyses in each of the hadronization models. As we have already mentioned in Sec. 6.2, for the dynamical model a proper matching to the parton shower (which allows for the shower cutoff $Q_{0}$ to be interpreted as an infrared factorization scale) implies a linear $Q_{0}$-dependence for the “hard” starting scales $\tilde{Q}_{g}$ and $\tilde{Q}_{q}$ of the non-perturbative branching processes and an insensitivity to $Q_{0}$ for the other hadronization parameters. On the other hand, for the default hadronization model (where the shower cutoff is merely another hadronization parameter) similar implications can in principle not necessarily be expected. In Figs. 8 and 9 we show the $Q_{0}$-dependence of the six tuning parameters which were treated as floating parameters for the default and the dynamic models, respectively, for $Q_{0}$ values between $1$ and $2$ GeV. We show the results based on the cubic (orange) and the quartic (blue) interpolations. The dots show the central values ${p_{i,\mathrm{cent}}(Q_{0})}$ and the vertical lines stand for the uncertainties which are obtained from the quadratic sum of the statistical und interpolation uncertainties $\Delta p_{i}=\sqrt{(\Delta p_{i,\mathrm{stat}}(Q_{0}))^{2}+(\Delta p_{i,\mathrm{inter}}(Q_{0}))^{2}}$, as described in Sec. 6.2. Overall we see that tuning results based on the cubic and quartic interpolations are nicely compatible, even though the error bars do not always overlap.
@@
In Fig. 9 we see that, indeed, the parameters $\tilde{Q}_{g}$ and $\tilde{Q}_{q}$ exhibit a $Q_{0}$-dependence that is quite close to linear particularly for $\tilde{Q}_{g}$, which governs the gluon splitting dynamics of the dynamical model. For $\tilde{Q}_{q}$ the expected behavior only sets in for $Q_{0}>1.3$ GeV, which is likely related to the fact that it governs the cluster fission dynamics that takes place at a later stage and may not have sufficient room to adapt for low $Q_{0}$ values. At the same time, in comparison the other four hadronization parameters (${\mathtt{Cl}_{\mathtt{max}}}$, ${\mathtt{Cl}_{\mathtt{pow}}}$, $\mathtt{PwtSquark}$ and $\mathtt{PwtDIquark}$) are, indeed, rather $Q_{0}$-insensitive. The behavior is again not perfect, with ${\mathtt{Cl}_{\mathtt{max}}}$ and ${\mathtt{Cl}_{\mathtt{pow}}}$ exhibiting rather large uncertainties, but it can nevertheless be observed rather clearly. In contrast, we can see a somewhat larger $Q_{0}$-dependence of the same four parameters for the default model in Fig. 8. This is visible most prominently for $\mathtt{PwtDIquark}$. Since these parameters directly affect the Baryon production and thus the charge particle multiplicities generated in the simulation, we can conclude that in the default model the shower cutoff value $Q_{0}$ affects these multiplicities so that the tuned value of $\mathtt{PwtDIquark}$ needs to compensate in a significant way. In the novel dynamical hadronization model this rather unnatural feature is not visible. We also note that the not completely smooth behavior visible for some of the model parameters as a function of $Q_{0}$ can be partly attributed to mass, threshold as well as some exceptional kinematic effects that can take place in the cluster model when $Q_{0}$ is varied. Overall, the results shown in Figs. 8 and 9 again confirm that the novel dynamical hadronization model achieves a better separation of the low-scale hadronization processes from the parton level description provided by the parton shower. This separation is essentially the MC simulation analogue of factorization in analytic QCD studies and an important prerequisite for the parton shower cutoff to be interpreted as a factorization scale.
@@ 7.3 Rescaled Migration Matrix Function
Before moving on, let us comment on a general feature of the migration functions. It concerns that $S^{\rm MC}(\ell,\{\hat{k},Q,Q_{0}\})$ is finite for $\ell$ values down to $-1$ GeV. We can see this feature clearly in the right panels of Figs. 10 and Fig. 11 for partonic values $\hat{k}>1$ GeV. Even though the major portion of $S^{\rm MC}(\ell,\{\hat{k},Q,Q_{0}\})$ is located at positive $\ell$, which means that the bulk of hadronization corrections shift the thrust distribution towards larger thrust values, this negative tail shows that hadronization can sometimes also decrease the thrust value. Since this feature arises from the tuning to actual LEP data, it means that at least for the Herwig simulation implementation this particular feature of the hadronization corrections is demanded by data. For $\hat{k}\geq 1.5$ GeV this behaviour arises consistently and becomes independent of the value of $\hat{k}$ also for other $Q$ and $Q_{0}$. This feature is per se not at all problematic and also consistent with a shape function within QCD factorization. However, for partonic $\hat{k}$ values below $1.5$ GeV the migration function is not capable to build up such a negative tail, simply because the physical thrust values are restricted to be positive. This entails that the first $\ell$-moment of the rescaled migration function $S^{\rm MC}(\ell,\{\hat{k},Q,Q_{0}\})$ (see Eq. (13)) for $\hat{k}$ values below $1.5$ GeV is always larger than for values larger than $2$ GeV, where the first moment stabilizes. This feature is absent in the shape function $S_{\rm had}(\ell,Q_{0})$, which is strictly $\hat{k}$-independent. We come back to this feature in our discussion below.
@@ 7.4 Shower Cutoff and Energy Dependence of the Migration Matrix Function
where $\Delta_{\rm soft}(Q_{0},Q_{0,\mathrm{ref}})$ is determined from the $R$-evolution equation in Eq. (10) employing the strong coupling extracted from Herwig in the $\overline{\rm MS}$ scheme.
@@
It is instructive to discuss this failure of the default model from a the perspective of overall size of the hadronization corrections. As we can see from Fig. 14, for $Q_{0,\mathrm{ref}}=1.25$ we have $\Omega_{1}^{\rm MC}$ in the range between $1.2$ and $1.5$ GeV for partonic momenta $\hat{k}>2$ GeV. This should be compared to the overall variation of $\Omega_{1}^{\rm MC}$ in the range between $Q_{0}=1.0$ GeV and $Q_{0}=1.75$ GeV predicted by $R$-evolution equation (10) obtained from QCD factorization. The latter amounts to around $0.5$ GeV, as we see in the lower panels of Fig. 15. However, because the choice of reference shower cut $Q_{0,\mathrm{ref}}$ is arbitrary, and we do not have any principle means to tell for which value of $Q_{0}$ the default model may have the best agreement with the expectations from QCD factorization (which can only be tested by a $Q_{0}$-dependence consistent with QCD factorization), this implies that the size of the hadronization corrections to the (true) partonic thrust distribution that is provided by the default model is inconsistent with QCD factorization in the range of $40$%. This is substantially worse in comparison to the new default model, where we have consistency at the level of better than $10$%. In the context of the shower cut $Q_{0}$ adopting the role of an IR factorization scale, so that the hadronization effects have a well-defined and controlled scheme dependence so that they can be assigned field theoretic meaning, the default model thus fails quite badly. In this respect the novel dynamical hadronization model performs substantially better.
@@ 8 Conclusions and Outlook
In this article we promote the idea of the parton shower cutoff $Q_{0}$ for MC simulations being an infrared factorization scale that separates, in a controlled manner and compatible with QCD, perturbative and non-perturbative hadronization effects. In this context, features of the MC’s hadronization model may be given a more systematic field theoretic meaning and, at the same time, QCD parameters appearing in the parton shower may be related to their renormalization scheme dependent counter parts appearing in analytic QCD computations. An important prerequisite to achieve this is, to have a parton shower algorithm that has at least NLL precision for the observable considered. However, the second essential and equally important prerequisite is to have a hadronization model that can properly match the unavoidable infrared cutoff $Q_{0}$-dependence of the parton level description that emerges from the parton shower. This is a highly nontrivial condition on the hadronization model since it entails that the parton shower cutoff $Q_{0}$ is not treated as a tuned hadronization parameter. Rather, the hadron level MC observable description should be equivalent for different $Q_{0}$ values, at least within some low energy interval, where QCD perturbation theory can still be trusted.
@@
In this work we have, for the first time in the literature, investigated the concept of the parton shower cutoff as a factorization scale coherently for a multipurpose MC event generator from the perspective of the interplay between the parton shower and the hadronization model, in this case for the coherent branching algorithm and the cluster model as implemented in Herwig 7. We have investigated these features using the angular ordered parton shower and the cluster hadronization model implemented in the Herwig 7.2 MC event generator focusing primarily on the 2-jettiness distribution in $e^{+}e^{-}$ annihilation for which the angular ordered parton shower is NLL precise. In the earlier work Hoang:2018zrp some of us have analyzed the gluon transverse momentum cutoff $Q_{0}$-dependence that emerges from the angular ordered parton shower for the 2-jettiness distribution in detail at the subleading ${\cal O}(\alpha_{s})$ level using QCD factorization in the dijet region. It was also shown that the Herwig parton shower implementation satisfies the $Q_{0}$ evolution equation from QCD factorization very well. In this work we have now extended our investigations concerning the parton shower infrared cutoff being a factorization scale from the perspective of the Herwig cluster hadronization model.
@@
The default cluster hadronization model has not been designed in a way so that it systematically matches to the parton shower for a range of $Q_{0}$ values. This is related to a number of features in the cluster formation and fission dynamics that yield good data description through the tuning, but are otherwise ad-hoc and not compatible with the processes that happen in the parton shower evolution close to the cutoff $Q_{0}$. As a result, the hadronization effects to the 2-jettiness distribution provided by the default Herwig cluster model do not satisfy the QCD constraints that emerge from $Q_{0}$ being promoted to a factorization scale in a satisfactory manner. We have demonstrated this in a number of tests based on tuning analyses for $Q_{0}$ shower cutoff values in the range between $1$ and $2$ GeV where $Q_{0}$ is treated as an external scale and not a tuned parameter. These tests involve analyses of the $Q_{0}$-dependence of hadron level simulations for 2-jettiness and also other event-shapes and observables and of the parton-to-hadron migration matrix for 2-jettiness for which QCD factorization provides nontrivial constraints.
@@
To improve the cluster hadronization model we have added a number of features to the cluster formation and cluster fission processes that better mimic the gluon emission and splitting dynamics that takes place in the parton shower. These modifications provide a clearer separation of model parameters that are expected to be correlated to $Q_{0}$, governing the ’hard’ aspects of the cluster hadronization dynamics, from those which govern ’soft’ hadron formation aspects that should be rather $Q_{0}$-independent. In our $Q_{0}$-dependent tuning studies we found that this novel dynamical cluster hadronization model performs substantially better concerning the $Q_{0}$-invariance of Herwig’s hadron level predictions as well as concerning the QCD factorization constraints on the 2-jettiness parton-to-hadron migration matrix. We emphasize that our analyses also involved hard scattering energies that are not accounted for in the reference data used for the tuning. This shows that the novel dynamical hadronization model properly scales the consistency to QCD factorization to other hard scattering energies. We have also investigated the hadron level simulation of many other $e^{+}e^{-}$ observables, including jet rates and multiplicities, and found globally a significant reduction of the $Q_{0}$-dependence in the dynamical model. However, since the concrete results for the parton level $Q_{0}$-dependence are currently only available for 2-jettiness and thrust related event-shape observables Hoang:2018zrp , a separate test of the parton level results and the hadronization corrections is not yet possible for other types of observables. To acquire such results a systematic knowledge of the infrared sensitivity of an observable (and thus the structure of its hadronization corrections) is mandatory.
@@
Even though the novel dynamical hadronization model we have designed in this article is not perfect, its features provide an important step forward in promoting the hadronization corrections encoded in MC generators to have a well-defined scheme in the QCD context, as also discussed previously in Ref. Platzer:2022jny . This is an essential aspect that should be followed in parallel to the ongoing developments of subleading order precise parton shower algorithms such that the parameters of the parton shower and eventually even of the hadronization model may acquire a systematic QCD field theoretic meaning. Beyond a more precise and consistent description of experimental data, important potential applications of such improvements are the determination of QCD parameters directly from MC studies as well as a systematic quantification of hadronization corrections to analytic QCD calculations from MC simulations. In an upcoming article we will investigate the former application from the perspective of the MC top quark mass parameter.